OpenStudy (itsme.caro):

find dy/dx. y=x2log2(3−2x)

jhonyy9 (jhonyy9):

x2 so here this 2 is exponent ? and the second 2 wan being the base of logarithm ?

OpenStudy (kasep10):

Have you tried chain rule?\[y=x^2 \log_2 (3-2x)\]

OpenStudy (itsme.caro):

i understand the chain rule, i get stuck in the log part, not sure how to derive this.

OpenStudy (3mar):

May I help?

OpenStudy (3mar):

Differentiation of logarithmic function. log and ln Feel free to discuss wherever you need

OpenStudy (itsme.caro):

ok let me see if i can figure it out

OpenStudy (3mar):

ok

OpenStudy (itsme.caro):

no i can't

OpenStudy (itsme.caro):

i have no idea where to go with this

OpenStudy (3mar):

Is this the whole function? \[y=x^2\log_2(3-2x)\]

OpenStudy (3mar):

@itsme.caro

OpenStudy (3mar):

Do you follow?

OpenStudy (itsme.caro):

what do you mean?

OpenStudy (3mar):

Is what I wrote correct?

OpenStudy (itsme.caro):

yes, that part is correct. i thought about using productivity rule, but i don't know how to differentiate log

OpenStudy (itsme.caro):

ok so so far i got this: 2xlogbase2 (3-2x) that part is right... but the second part of the productivity rule is f g' which would be x^2 (this part i don't know)

OpenStudy (3mar):

productivity rule \[\frac{ d }{ dx }[f(x).g(x)]=f(x).g'(x)+f'(x).g(x)\] So it will be like that: \[=(x^2).[(\log_2e).\frac{ 1 }{ 3-2x }]+\log_2(3-2x).[2x]\] I think it is clear now.

OpenStudy (3mar):

Did you get it? @itsme.caro

OpenStudy (itsme.caro):

so when you get the derivative the e shows up?

OpenStudy (karim728):